40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jul 23 15:00:28 2017, Jamie, Omnistructure, VAC, strange sound around X arm vacuum pumps 
    Reply  Mon Jul 24 09:55:29 2017, Steve, Omnistructure, VAC, controller failer TP2controllerFails.png
       Reply  Tue Jul 25 08:48:57 2017, Steve, Update, VAC, RGA scan at d278 RGA278d.png
          Reply  Mon Aug 28 10:28:07 2017, Steve, Update, VAC, RGA scan 20 day after 17 Torr 20d_after_17_torr.png
             Reply  Tue Jan 30 10:22:20 2018, Steve, Update, VAC, RGA scan at day 175 RGApd80-b-468-175d.pngRgaOn20hrs.png
                Reply  Fri Mar 16 09:36:56 2018, Steve, Update, VAC, RGA scan at day 511,218d RgaScan511d.png08.png
                   Reply  Mon Mar 26 09:25:18 2018, Steve, Update, VAC, VM1 opened CC4VM1.png
                      Reply  Tue Mar 27 08:58:21 2018, Steve, Update, VAC, VM1 opened.......scan fine rga2018march27.png
          Reply  Tue Sep 19 11:05:26 2017, Steve, Update, VAC, RGA scan at day 334  RGAscan334d.png
             Reply  Fri Oct 27 10:14:11 2017, Steve, Update, VAC, RGA scan at day 372 RGA_scan_d372.png
                Reply  Mon Oct 30 10:09:41 2017, Steve, Update, VAC, TP2 failed tp2failed.pngifo_1.0E-5_Torrit.pngtp2failed2dago.png4days.png
                   Reply  Mon Nov 6 09:09:43 2017, Steve, Update, VAC, setting up new TP2 turbo 
       Reply  Tue Jul 25 14:04:06 2017, Steve, Update, VAC, turbo controller installed and we are running at vac normal TP2@75Krpm.png
Message ID: 13260     Entry time: Mon Aug 28 10:28:07 2017     In reply to: 13142     Reply to this: 13588
Author: Steve 
Type: Update 
Category: VAC 
Subject: RGA scan 20 day after 17 Torr 

The RGA was turned on 7 days ago.  It's 46 C now.   The X-arm room tem ~20 C

IFO pressure 6.5e-6 Torr at IT-Hornet gauge. Valve configuration vacuum normal.

Attachment 1: 20d_after_17_torr.png  44 kB  | Hide | Hide all
20d_after_17_torr.png
ELOG V3.1.3-